1/1

Werner Bergengruen,

Ein Vorfrühlingstag

Argiope information: The view of this version is restricted. Presumably there are copyright issues.

Author of this German poem is Werner Bergengruen (∗1892-09-16 – †1964-09-04).

Arg-467-887 (2008-02-21 11:20:49)

http://www.poezio.net/version?poem-id=467&version-id=887